OpenStudy (anonymous):

Jackie has a jar of 54 coins consisting of dimes and quarters. If she has a total of $10.05, how many dimes does she have?

OpenStudy (anonymous):

uhh..23 maybe?

OpenStudy (anonymous):

wait no..

OpenStudy (turingtest):

how are you trying to set up the problem?

OpenStudy (anonymous):

31?

OpenStudy (turingtest):

or were you just guessing?

OpenStudy (anonymous):

messing around with it in my graphing calculator

OpenStudy (anonymous):

ive done several different things and gotten 23 and 31

OpenStudy (turingtest):

well I can't help you with your graphing calculator I can help you with the math though, which is the important part.

OpenStudy (anonymous):

alright im listening

OpenStudy (anonymous):

help please

OpenStudy (turingtest):

you can make a set of two equations out of this let x=#dimes let y=#quarters

OpenStudy (turingtest):

"Jackie has a jar of 54 coins consisting of dimes and quarters" we can interpret this mathematically as\[x+y=54\]making sense?

OpenStudy (anonymous):

yes so far

OpenStudy (turingtest):

now can you come up with an expression for how much money there is total in terms of x and y ?

OpenStudy (anonymous):

x+y=$10.05

OpenStudy (turingtest):

not quite... You will need coefficients for this one that relate to the value of each coin. For instance, if I have x quarters, then the amount of money I have in quarters is 0.25x (in dollars) So here we have quarters \(and\) dimes adding to a total of $10.05 So what will the expression be?

OpenStudy (anonymous):

.25x+.10y=10.05

OpenStudy (turingtest):

good :)

OpenStudy (anonymous):

this isnt so much what i had trouble with

OpenStudy (anonymous):

i tried several different formulas and got several answers which i think are right lol.

OpenStudy (anonymous):

it would help me alot if you could explain the use of the correct formula

OpenStudy (turingtest):

the point is that now we have two equations with two variables, which can be solved in a number of ways \[x+y=54\]\[0.25x+0.1y=10.05\]I recommend the first step is to get rid of the decimals in the second equation can you do that?

OpenStudy (anonymous):

normally to get rid of a variable you divide. its been awhile since i did this i forgot mutliple variables

OpenStudy (turingtest):

I didn't say get rid of a variable, I said get rid or the decimals in the second equation

OpenStudy (anonymous):

i apreciate the help by the way :)

OpenStudy (turingtest):

you're welcome :)

OpenStudy (anonymous):

um..

OpenStudy (anonymous):

-.25?

OpenStudy (anonymous):

and .10?

OpenStudy (anonymous):

that seems ridiculous to me

OpenStudy (turingtest):

what would happen if you multiplied both sides by 100 in the second equation?

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

haha

OpenStudy (turingtest):

right, now it get's pretty easy pretty fast

OpenStudy (anonymous):

totally forgot about doing that

OpenStudy (anonymous):

yea lol

OpenStudy (turingtest):

ok, so what is the second equation after multiplying by 100 ?

OpenStudy (anonymous):

25x +10y = 10.05

OpenStudy (turingtest):

not exactly I think you forgot to multiply \(both\) sides

OpenStudy (anonymous):

1005 :)

OpenStudy (anonymous):

= 1005 soryr

OpenStudy (anonymous):

does moderator mean your employed by this site or just an admin?

OpenStudy (turingtest):

right, so now our system is\[x+y=54\]\[25x+10y=1005\]now I would prescribe the substitution method from here

OpenStudy (turingtest):

I'm just a moderator, which basically means I deal with immature trolls and such I also try to contribute ideas to administration

OpenStudy (anonymous):

cool, did you start out as a user ?

OpenStudy (turingtest):

yup, only been a mod for a couple weeks actually. anyhow... do you know how to solve systems?

OpenStudy (anonymous):

ya one sec, i think i have the anser

OpenStudy (turingtest):

you used a computer lol

OpenStudy (turingtest):

..or are you doing it by hand?

OpenStudy (anonymous):

hand

OpenStudy (anonymous):

i messed up tho

OpenStudy (turingtest):

good :) that is a consequence, but the practice is important

OpenStudy (anonymous):

lol

OpenStudy (turingtest):

it's easy to mess up with systems like this before you learn linear algebra anyhow, what sub did you make? can you show me what you did?

OpenStudy (anonymous):

linear algebra?

OpenStudy (anonymous):

is that tought in algebra 1 or 2

OpenStudy (turingtest):

nevermind... higher math

OpenStudy (anonymous):

its actually sad but ive taken agebra 1, 2, advanced with financial app and advanced geo

OpenStudy (turingtest):

no it deals with matrices and such... it's usually around when you take calculus

OpenStudy (anonymous):

the worst part is i slacked hard and didnt bother to atcually LEARN the work...just got as many correct asnwers asap

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

anyway

OpenStudy (anonymous):

whats the next step

OpenStudy (anonymous):

o wait

OpenStudy (anonymous):

can you do it normally?

OpenStudy (anonymous):

-10y?

OpenStudy (anonymous):

etc?

OpenStudy (anonymous):

25x + 10y + -10y = 1005 + -10y

OpenStudy (anonymous):

if you did that couldnt you combine like terms and then solve?

OpenStudy (turingtest):

yeah, we used to contribute to that they made me a mod partially to stop that habit and start teaching more we have equation that we can name\[1)x+y=54\]\[2)25x+10y=1005\]and what you are doing would work if your "etc." is right, but you are doing it the hard way you cannot combine like terms and solve yet because you still have both an x and a y

OpenStudy (anonymous):

x = 40.2 - 0.4y thats the answer but i dont know how to convert to just one number

OpenStudy (anonymous):

i think?

OpenStudy (anonymous):

if not that way then i dont know

OpenStudy (turingtest):

I'm showing you, please try to be patient I'm typing it out

OpenStudy (anonymous):

and yea, open study used to be alot more pure answers instead of breaking it down

OpenStudy (anonymous):

sorry, i was just elaborating lol

OpenStudy (turingtest):

\[1)x+y=54\]\[2)25x+10y=1005\]we can turn equation 1) around and solve it for y\[y=54-x\]now we can substitute this expression in for \(y\) in equation 2) and it will only contain x's, which means we can then solve for x what do you get by subbing \(y=54-x\) into equation 2) ???

OpenStudy (anonymous):

25x+540=1005

OpenStudy (anonymous):

1005 - 540 = 456

OpenStudy (anonymous):

465*

OpenStudy (anonymous):

thank you :)

OpenStudy (turingtest):

? you lost art of the expression there should be another x term

OpenStudy (turingtest):

*part of the...

OpenStudy (anonymous):

hmm?

OpenStudy (anonymous):

i got 18

OpenStudy (turingtest):

\(y=54-x\) subbing that into \[25x+10y=1005\]gives\[25x+10(54-x)=1005\]

OpenStudy (anonymous):

let number of dimes be x therefore quarters would be 54-x 0.1x + 0.25(54-x) = 10.05 now solve for x

OpenStudy (turingtest):

what x= I don't yet know, I haven't checked

OpenStudy (turingtest):

I am just prescribing the approach

OpenStudy (anonymous):

now im confused again lol

OpenStudy (anonymous):

why would you add another x into the equation

OpenStudy (turingtest):

@sheg yes you are right, I was going to do elimination so I multiplied the other one first

OpenStudy (turingtest):

though both give the same answer of course

OpenStudy (turingtest):

@rkasser that is what we got by solving equaiton 1) for y as I explained above

OpenStudy (anonymous):

hey rkasser total number of coins = 54 and if out of 54 coins x are dimes so how many are left??/ it should be somewhat 54-x

OpenStudy (turingtest):

@sheg I have said this already, please read above

OpenStudy (anonymous):

So number of dimes = x number of quarters = 54-x